OpenStudy (anonymous):

solve: sqrt x+4=-10 and sqrt x+4=10. How do I do this?

hartnn (hartnn):

i would say , square both sides what u get ?

OpenStudy (anonymous):

separate questions?

hartnn (hartnn):

i think he is trying to solve \(\large |\sqrt{x+4}|=10\)

OpenStudy (anonymous):

oooh

hartnn (hartnn):

or she , if she is a girl

OpenStudy (anonymous):

im a girl. and they are two separate questions. and I don't speak algebra, lol. I am not good at this -__-

hartnn (hartnn):

oh lol :P anyways square both sides, what u get ?

OpenStudy (anonymous):

\[\sqrt{x+4}=-10 \] for first one?

OpenStudy (anonymous):

Yeah square both sides, then solve for x.

OpenStudy (anonymous):

4 squared would be 16, right? and 10 squared would be 100? am I on the right track??

OpenStudy (anonymous):

Why did you square the 4?

OpenStudy (anonymous):

\[(\sqrt{x+4})^2=(-10)^2\]

OpenStudy (anonymous):

lol, idk. I am so confused. :/

OpenStudy (anonymous):

x=6?

OpenStudy (anonymous):

no, wait... x=96? ughhh. I am confused!

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

now do the same for the other one

OpenStudy (anonymous):

so x=96 or x=6?

OpenStudy (anonymous):

96 idk where you got the 6 from

OpenStudy (anonymous):

lol, I didn't square it at first. umm, for the other problem.. wouldn't it still be 96?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

If I have a problem:  and my instructions are to multiply and simplify. assume that all variables represent nonnegative numbers. how would I do that?